Let c0 be the space of sequences of complex numbers which converge to 0. That is

c_0 = {(x_i)_i∈N : xi ∈ C, x_i → 0}.

(i) Show that c0 is a closed subspace of L^∞.

(ii) Define a mapping T by

T : L^∞ → c_0

(x_n)_n → (x_n/ n ) _n

Show that T is a (linear bounded) operator. Show that ran T is not closed.


1
Expert's answer
2021-02-24T06:10:03-0500

i) Let "x^{(k)}\isin c_0" be a sequence converging to "w\isin L^{\infin}". Take "\epsilon>0" and "N_0\isin N" such that

"\displaystyle \sup_{1\leq i\leq \infin}|x^{(k)}_i-w_i|<\frac{\epsilon}{2}"

for all "k>N_0". For each "k" choose "N_1\isin N" such that

"|x^{(k)}_i|\leq \frac{\epsilon}{2}" for all "i>N_1" .

Thus, "|w_i|=|w_i-0|\leq |w_i-x^{(k)}_i|+|x^{(k)}_i|\leq \epsilon" for "i>N_1", and "k>N_0" , that means that the sequence "w_i" converges to and "w_i\isin c_0" . Hence "c_0" is closed in "L^{\infin}" .


ii) For linear bounded) operator "T:V_1\to V_2"

"||Tv||_2\leq C||v||_1," constant "C>0"

In our case:

for "L^{\infin}: ||x||=sup|x_n|"

for "c_0" : "||x||=sup|x_n\/n|"


Then:

"sup|x_n/n|\leq Csup|x_n|"

So, there exists "C>0" . That is, "T" is a linear bounded operator.


"Ran(T)=(x_n\/n)_n"

If sequence "(x_n)" in "L^{\infin}" does not converge, then sequence "(x_n\/n)_n" does not converge to 0.

So, "Ran(T)" is not closed.